A wire is sitting flat on a page, with a constant current moving north (top of the page.) The wire is between the poles of a magnet, with the north pole to the right of the wire, and the south pole to the left. What is the direction of the magnetic force on the current carrying wire A. Into the page . B. Toward the top of the page. C. Toward the bottom of the page. D. out of the page

A negative charge q moves toward the left on the paper you are looking at with a velocity v. A constant magnetic field points directly out of the page. What is the direction of the magnetic force on the charge? A. Toward the top of the page. B. Toward the right. C. Toward the left. D. Toward the bottom of the page.
Answer #1

The current in the wire is in north direction, therefore the electrons inside the wire are moving in the south direction.

Force on the electron is given by

F=q (\vec v\times\vec B) where \vec v is speed of electrons.

F=-e(\vec v\times\vec B)

This tells that the force is in opposite direction to (\vec v\times\vec B) .

The direction of (\vec v\times\vec B) can be found using Right Hand Thumb rule. Point your index finger in the direction of velocity, middle finger in the direction of magnetic field. Your thumb will give the direction of (\vec v\times \vec B) . The direction of (\vec v\times \vec B) is into the page.

Therefore, the direction of magnetic force on every electron inside the wire is out of the page opposite to the direction of (\vec v\times\vec B) .

Hence, the direction of magnetic force on the current carrying wire is Out of the page.

(D) Out of the page is the correct option.

(2) The magnetic force on the charge q is

F=q (\vec v\times \vec B)

Again, it is given that q is negative, therefore the direction of the magnetic force on the charge is in opposite direction to (\vec v\times \vec B) .

The direction of (\vec v\times \vec B) can be found using Right Hand Thumb rule.

Point your index finger in the direction of velocity, your middle finger in the direction of magnetic field. Your thumb will give the direction of (\vec v\times \vec B) . The direction of (\vec v\times \vec B) is top of the page.

The direction of the magnetic force on negative charge q is Bottom of the page opposite to the direction of (\vec v\times \vec B) .

(D) towards bottom of the page is the correct option.